OpenStudy (anonymous):

Help? :) What is the slope and y-intercept of this line? y = –6x + 2 A. slope: –6, y-intercept: 2 B. slope: –3, y-intercept: 1 C. slope: 2, y-intercept: –6 D. slope: 6, y-intercept: –2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks =D

OpenStudy (anonymous):

can you help with acouple more?

OpenStudy (anonymous):

sure i can try

OpenStudy (anonymous):

ok thanks ;) , What is the slope-intercept form of this equation? 3x + 9y = 18 A. y=(-1/3)x+6 B. y=(-1/3)x+2 C. y=1/3x+6 D. y=2/3x+2

OpenStudy (anonymous):

put it into y=mx+b form where m is the slope and b is the y intercept

OpenStudy (anonymous):

so 9y=-3x+18 divide by 9 y=-1/3x+2

OpenStudy (anonymous):

awesome! XD Im really bad when it come's to these things but can you help me out with 1 last one?

OpenStudy (anonymous):

sure :3

OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = x – 5 A. slope: –5, y-intercept: 1 B. slope: –1, y-intercept: 5 C. slope: 5, y-intercept: –1 D. slope: 1, y-intercept: –5

OpenStudy (anonymous):

D ;)

OpenStudy (anonymous):

omg... your a life saver! XD thank you soo much

OpenStudy (anonymous):

np ;3

OpenStudy (anonymous):

have a great day :) XD

OpenStudy (anonymous):

you too :3